# Valentina Bais and Rafael Torres Scuola Internazionale Superiori di Studi Avanzati (SISSA) Via Bonomea 265 34136

### v1

@@ 3. III
The crux to the efficiency of the procedure is its fourth step. In principle, one would hope that the diffeomorphism invariant that discerns $M_{1}$ from $M_{2}$ also tells apart the new four-manifolds $M_{1}(p)$ and $M_{2}(p)$. This is indeed the case in all the examples under consideration in this paper, where we reckoned the Heegaard Floer homology based invariants used by Levine-Lidman-Piccirillo in [27], the Ozsváth-Szabó invariants [29], the Seiberg-Witten invariants [19], [31, Section 2.4] (as employed by Stipsicz-Szabó [30, 31] and Baykur-Stipsicz-Szabó [8]), and the $\eta$-invariant as studied by Stolz [32].
@@
To distinguish the smooth structures in [27, Theorem 1.2], Levine-Lidman-Piccirillo show that the universal covers ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{E}$ of the four-manifolds $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{R}$ are non-diffeomorphic by means of their Ozsváth-Szabó invariant [29]. For every odd $p\in\mathbb{Z}_{>0}$, we show that the degree-two coverings of our four-manifolds are obtained from ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{E}$, respectively, by taking a connected sum with a $\mathbb{Q}$-homology four-sphere. This observation, together with the properties of the Heegaard Floer smooth four-dimensional invariant, allows us to canonically adapt the proof of [27, Theorem 1.2] to distinguish such coverings and, hence, also the four-manifolds $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{R}_{2p}$.
Our next theorem was inspired by recent work of Stipsicz-Szabó [30, 31] and Baykur-Stipsicz-Szabó [8], where they built infinitely many nondiffeomorphic irreducible smooth structures on $\Sigma_{4m+2}\mathbin{\#}\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ for $\Sigma_{4m+2}$ a $\mathbb{Q}$-homology four-sphere with fundamental group $\mathbb{Z}/(4m+2)$. Recall that a smooth four-manifold $M$ is said to be irreducible if for every smooth connected sum decomposition $M=M_{1}\mathbin{\#}M_{2}$ either $M_{1}$ or $M_{2}$ is a homotopy four-sphere [19, Definition 10.1.17].
@@
    $\tau_{n}:\widetilde{X}_{n}\rightarrow\widetilde{X}_{n}$
@@
    $X_{n}(2p)\rightarrow Y\mathbin{\#}\Sigma_{2p,k}.$
@@
    $\widehat{X}_{n}(p)=\widetilde{X}_{n}\mathbin{\#}\Sigma_{p}.$
@@
In [30, 31, 8], the Seiberg-Witten invariants are used to distinguish the diffeomorphism types of their examples in the following way. Notice that a diffeomorphism $h:M_{1}\rightarrow M_{2}$ between two closed oriented four-manifolds lifts to a diffeomorphism of the degree-two coverings $\hat{h}:\widehat{M}_{1}\rightarrow\widehat{M}_{2}$, whenever the index-two subgroups of $\pi_{1}(M_{i})$ identifying $\widehat{M}_{i}\rightarrow M_{i}$ correspond under the map induced by $h$; see Lemma 8. Stipsicz-Szabó obstruct the existence of the diffeomorphism $h$ through the diffeomorphism invariance of the Seiberg-Witten function $\operatorname{SW}_{\widehat{M}_{i}}:\operatorname{Spin}^{\mathbb{C}}(\widehat{M}_{i})\rightarrow\mathbb{Z}$ for $i=1,2$. Moreover, they conclude on the irreducibility of $\widehat{M}_{i}$ and $M_{i}$ through their Seiberg-Witten basic classes. We follow the same approach in the proof of Theorem D. In our case, however, the degree-two coverings of the four-manifolds of Theorem D do decompose smoothly as a connected sum $M\mathbin{\#}\Sigma_{p}$. In the proof of Theorem D, we couple the computations of the Seiberg-Witten invariants of Stipsicz-Szabó with a result of Kotschick-Morgan-Taubes [25, Proposition 2] that states that there are $\text{Spin}^{\mathbb{C}}$-structures on $M$ and $M\mathbin{\#}\Sigma_{p}$ for which the values of their Seiberg-Witten invariants coincide; see Section 4.1.
@@
The case of $m\neq 0\mod 2$ and $m\leq n$ of Theorem E is well-known to hold for every finite cyclic fundamental group and for infinitely many pairwise nondiffeomorphic irreducible smooth structures by modifying the constructions of closed simply connected irreducible four-manifolds of Akhmedov-Baldridge-Baykur-Kirk-Park [2], Akhmedov-Park [3], Baldridge-Kirk [5, 6], Baykur-Hamada [7], Fintushel-Stern [13] and Gompf [17] as already done in [3, 6, 8, 34, 35]. The case $m=0\mod 2$ and $p=1$ is due to Stipsicz-Szabó [30, 31], Baykur-Stipsicz-Szabó [8] and Arabadji-Porter [4]. Our proof of Theorem E is based on coupling their results with Theorem D, and it is presented in Section 5.
Concerning the nonorientable realm, we construct new smooth structures on $\mathbb{Q}$-homology real projective four-spaces through our mechanism, and extend results of Cappell-Shaneson [10] and Fintushel-Stern [11]. In order to build the novel four-manifolds, we start with Cappell-Shaneson’s exotic $\mathbb{R}P^{4}$. We address the fourth step of our construction mechanism by computing the $\eta$-invariant [32] of the four-manifolds involved in Section 6.4. Our main result in this case is the following theorem.
@@ 1.1. Acknowledgements
V. B. is a member of GNSAGA – Istituto Nazionale di Alta Matematica ‘Francesco Severi’, Italy. R. T. thanks Bob Gompf, Paul Kirk, Tye Lidman, Mark Powell and Zoltan Szabó for useful e-mail exchanges.
@@ 2.1. Cut-and-paste procedure: oriented setting
where $r_{\theta}$ is the rotation of $S^{2}$ of angle $\theta$ around a fixed axis [16], [19, p. 156]. Such map is called Gluck twist. In order to prove Theorem 4, it is enough to show that the diffeomorphism $\varphi$ admits an extension $\Phi_{p}:S_{p}\rightarrow S_{p}$ for any odd $p\in\mathbb{Z}_{>0}$. We explicitly build $\Phi_{p}$ as follows, using the identification
@@
coming from Lemma 2. Without loss of generality, we work in the setting of Remark 3. In particular, the deck transformation $\phi$ and the collection of three-balls $B_{1},\dots,B_{p}\subset D^{2}\times S^{1}$ are defined as in the proof of Lemma 2.
@@
where $r_{t}$ denotes the rotation of $S^{3}$ of an angle $2\pi t$ that restricts to a rotation along the $D^{2}$-factor on the solid torus $D^{2}\times S^{1}$ of the Heegaard decomposition of $S^{3}$ in Remark 3. Such solid torus contains the three-balls $B_{1},\dots,B_{p}$. Notice that for every $t\in[0,1]$, the map $F(t,\cdot)$ is an $f_{p}$-equivariant self-diffeomorphism of $S^{3}\setminus\sqcup_{i=1}^{p}B_{i}$, where $f_{p}$ can be chosen to act on the genus one Heegaard splitting of $S^{3}$ described in Remark 3.
We can now define the automorphism
@@
As mentioned in the first step of our introduction, remove the tubular neighborhood of a framed simple loop $\alpha_{M}\subset M$ from a closed oriented four-manifold $M$ to obtain a compact oriented four-manifold $M\setminus\nu(\alpha_{M})$ with boundary $\partial(M\setminus\nu(\alpha_{M}))=S^{1}\times S^{2}$. Cap off its boundary with the compact four-manifolds defined in (2.1) to construct a closed oriented four-manifold
@@
The reader is directed to [15, Chapter 5] regarding existence and uniqueness results of the tubular neighborhood of a locally flat simple loop in a topological four-manifold.
Independently of Corollary 5, the simplicity of the construction makes it straight-forward to compute several invariants of the new four-manifold (2.5), as we now sample.
@@
Let $\alpha_{M}\subset M$ be a simple loop whose homotopy class is $[\alpha_{M}]=0\in\pi_{1}(M)$. There is a diffeomorphism
@@ 3.1. Raw materials for Theorem C
We will employ as a raw material the exotic four-manifolds $\mathcal{E}$ and $\mathcal{R}$ constructed in [27, Theorem 1.2].
@@
We refer the reader to [27, Section 6.2] for the precise definition of the four-manifolds $\mathcal{E}$ and $\mathcal{R}$. The four-manifolds of Theorem C are defined by applying the construction described in Section 2.1 to the exotic four-manifold $\mathcal{R}$ in the second item of Theorem 9.
@@
Alternatively, the existence of a homeomorphism as in Lemma 10 follows from work of Hambleton-Kreck [20, Theorem C]; cf. [31, Theorem 2.1].
@@
[Proof]This follows from the definition of $\mathcal{R}$ in [27] as the quotient of $\mathcal{E}$ by a fixed-point-free orientation-reversing involution, together with Lemma 2 and Lemma 7.
@@ 3.2. Diffeomorphism obstruction
In light of the definition of $\mathcal{E}$ in [27, Section 6.2], Proposition 12 implies the existence of a decomposition
@@
where $V=C_{0}\cup_{\phi_{0}\circ\sigma}Z$ is the building block defined in [27, Construction 6.4]. In particular, $C_{0}$ is a contractible four-manifold, $Z$ consists of five $-1$-framed $2$-handles attached to $\partial C_{0}$, $\partial V$ is the result of 00-surgery on $S^{3}$ along the Stevedore knot and $\zeta$ is any of its orientation-reversing fixed-point-free involutions. On the other hand, for every $p\geq 1$, $V_{p}$ is defined to be the four-manifold
@@
where $V^{\prime}=C_{0}\cup_{\phi_{0}}Z$ and $V^{\prime}_{p}=C_{0}^{p}\cup_{\phi_{0}}Z$ are made up by the same building blocks as $V$ and $V_{p}$ with different gluing maps. We refer the reader to [27, Construction 6.4] for further details.
@@
We proceed to provide some technical lemmas that are needed to distinguish the smooth structures of Theorem C. These results are generalizations of [27, Proposition 6.9] and [27, Lemma 6.3] respectively. In particular, we make use of the Ozsváth-Szabó four-manifold invariants defined in [29, Section 2.4].
In the following, we denote by $\Psi_{W,\mathfrak{t}}\in HF^{-}_{\text{red}}(\partial W,\mathfrak{t}|_{\partial W})$ the relative invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(W,\mathfrak{t})$ where $W$ is a compact four-manifold with non-empty boundary, and $\Phi_{X,\mathfrak{s},L}$ is the closed four-manifold invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s})$ with respect to a one-dimensional subspace $L\subset H_{2}(X;\mathbb{Q})$. We refer the reader to [27, Section 6.1] for the needed background on such invariants.
@@
[Proof]The case $p=1$ is [27, Proposition 6.9]. Since $C_{0}^{p}\setminus B^{4}$ is a rational homology cobordism between $S^{3}$ and $\partial C_{0}^{p}$, the same proof holds for $p>1$ by replacing $C_{0}$ with $C_{0}^{p}$ in the argument. The proof of the following result is a straightforward adaptation of the one of [27, Lemma 6.3].
@@
[Proof of Proposition 13] Item (1) follows from the construction and the fact that $\mathcal{E}^{\prime}$ is an exotic ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ by [27, Theorem 6.7]. By Lemma 7, there is a degree-two covering
@@
Let us now prove Item (2). In particular, $\mathcal{E}_{p}$ and $\mathcal{E}^{\prime}_{p}$ are homeomorphic by construction, since $\mathcal{E}$ and $\mathcal{E}^{\prime}$ are homeomorphic by [27, Theorem 6.7]. To conclude the proof, we need to show that the degree-two coverings $\mathcal{E}_{p}$ and $\mathcal{E}^{\prime}_{p}$ of $\mathcal{R}_{2p}$ and $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ respectively are not diffeomorphic. Consider the decomposition
@@
where $V_{a}$ and $V_{b}^{p}$ are copies of $V$ and $V_{p}$ respectively without the $T^{-}_{Q_{0}}$ component, see [27, Construction 6.4]. As for the case of $\mathcal{E}$, we see that $H_{2}(\mathcal{E}_{p})$ has a basis
@@
  • $(\Xi,\Theta)$ is the basis for $H_{2}(T_{Q_{0}})$ from [27, Remark 3.11]. In particular, $\Xi$ is represented by a capped-off Seifert surface for $Q_{0}$ in $S^{3}(Q_{0})$ and $\Theta$ is represented by a $2$-sphere of self-intersection $-2$ and $\Xi\cdot\Theta=1$.
As a consequence of Lemma 14 and 15, there is a ${\operatorname{Spin}}^{\mathbb{C}}$-structure $(\mathcal{E}_{p},\mathfrak{t})$ on such that $\Phi_{\mathcal{E}_{p},\mathfrak{t},L}\neq 0$, where $L=\text{Span}(\Xi)\subset H_{2}(\mathcal{E}_{p};\mathbb{Q})$. By [27, Lemma 6.2], we have that $\Xi$ cannot be represented by any surface of genus less than $2$. It is possible to show that if there was a diffeomorphism
@@
then one could build a spherical representative for $\Xi$ as in the proof of [27, Theorem 6.7], which would lead to a contradiction.

### v2

@@ 3. III
The crux to the efficiency of the procedure is its fourth step. In principle, one would hope that the diffeomorphism invariant that discerns $M_{1}$ from $M_{2}$ also tells apart the new four-manifolds $M_{1}(p)$ and $M_{2}(p)$. This is indeed the case in all the examples under consideration in this paper, where we reckoned the Heegaard Floer homology based invariants used by Levine-Lidman-Piccirillo in [29], the Ozsváth-Szabó invariants [31], the Seiberg-Witten invariants [20], [34, Section 2.4] (as employed by Stipsicz-Szabó [33, 34] and Baykur-Stipsicz-Szabó [8]), and the $\eta$-invariant as studied by Stolz [35].
@@
To distinguish the smooth structures in [29, Theorem 1.2], Levine-Lidman-Piccirillo show that the universal covers ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{E}$ of the four-manifolds $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{R}$ are non-diffeomorphic by means of their Ozsváth-Szabó invariant [31]. For every odd $p\in\mathbb{Z}_{>0}$, we show that the degree-two coverings of our four-manifolds are obtained from ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{E}$, respectively, by taking a connected sum with a $\mathbb{Q}$-homology four-sphere. This observation, together with the properties of the Heegaard Floer smooth four-dimensional invariant, allows us to canonically adapt the proof of [29, Theorem 1.2] to distinguish such coverings and, hence, also the four-manifolds $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{R}_{2p}$.
Our next theorem was inspired by recent work of Stipsicz-Szabó [33, 34] and Baykur-Stipsicz-Szabó [8], where they built infinitely many nondiffeomorphic irreducible smooth structures on $\Sigma_{4m+2}\mathbin{\#}\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ for $\Sigma_{4m+2}$ a $\mathbb{Q}$-homology four-sphere with fundamental group $\mathbb{Z}/(4m+2)$. Recall that a smooth four-manifold $M$ is said to be irreducible if for every smooth connected sum decomposition $M=M_{1}\mathbin{\#}M_{2}$ either $M_{1}$ or $M_{2}$ is a homotopy four-sphere [20, Definition 10.1.17].
@@
    (1.1) $\tau_{n}:\widetilde{X}_{n}\rightarrow\widetilde{X}_{n}$
@@
    (1.2) $X_{n}(2p)\rightarrow Y\mathbin{\#}\Sigma_{2p,k}.$
@@
    (1.3) $\widehat{X}_{n}(p)=\widetilde{X}_{n}\mathbin{\#}\Sigma_{p}.$
@@
If the four-manifold $\widetilde{X}_{n}$ does admit an involution (1.1), but has more than two Seiberg-Witten basic classes, assume it is minimal. Then, there exists a closed smooth oriented irreducible four-manifold $X_{n}(2p)$ with finite fundamental group of order $2p$ for which there is a homeomorphism (1.2) whose double cover is (1.3).
In [33, 34, 8], the Seiberg-Witten invariants are used to distinguish the diffeomorphism types of their examples in the following way. Notice that a diffeomorphism $h:M_{1}\rightarrow M_{2}$ between two closed oriented four-manifolds lifts to a diffeomorphism of the degree-two coverings $\hat{h}:\widehat{M}_{1}\rightarrow\widehat{M}_{2}$, whenever the index-two subgroups of $\pi_{1}(M_{i})$ identifying $\widehat{M}_{i}\rightarrow M_{i}$ correspond under the map induced by $h$; see Lemma 9. Stipsicz-Szabó obstruct the existence of the diffeomorphism $h$ through the diffeomorphism invariance of the Seiberg-Witten function $\operatorname{SW}_{\widehat{M}_{i}}:\operatorname{Spin}^{\mathbb{C}}(\widehat{M}_{i})\rightarrow\mathbb{Z}$ for $i=1,2$. Moreover, they conclude on the irreducibility of $\widehat{M}_{i}$ and $M_{i}$ through their Seiberg-Witten basic classes. We follow the same approach in the proof of Theorem D. In our case, however, the degree-two coverings of the four-manifolds of Theorem D do decompose smoothly as a connected sum $M\mathbin{\#}\Sigma_{p}$. In the proof of Theorem D, we couple the computations of the Seiberg-Witten invariants of Stipsicz-Szabó with a result of Kotschick-Morgan-Taubes [27, Proposition 2] that states that there are $\text{Spin}^{\mathbb{C}}$-structures on $M$ and $M\mathbin{\#}\Sigma_{p}$ for which the values of their Seiberg-Witten invariants coincide; see Section 4.1.
@@
The case of $m\neq 0\mod 2$ and $m\leq n$ of Theorem E is well-known to hold for every finite cyclic fundamental group and for infinitely many pairwise nondiffeomorphic irreducible smooth structures by modifying the constructions of closed simply connected irreducible four-manifolds of Akhmedov-Baldridge-Baykur-Kirk-Park [2], Akhmedov-Park [3], Baldridge-Kirk [5, 6], Baykur-Hamada [7], Fintushel-Stern [14] and Gompf [18] as already done in [3, 6, 8, 37, 38]. The case $m=0\mod 2$ and $p=1$ is due to Stipsicz-Szabó [33, 34], Baykur-Stipsicz-Szabó [8] and Arabadji-Morgan [4]. Our proof of Theorem E is based on coupling their results with Theorem D, and it is presented in Section 5.
Concerning the nonorientable realm, we construct new smooth structures on $\mathbb{Q}$-homology real projective four-spaces and non-smoothable copies of these four-manifolds through our mechanism, and extend results of Cappell-Shaneson [10], Fintushel-Stern [12] and Ruberman [32]. In order to build the novel four-manifolds, we start with Cappell-Shaneson’s exotic $\mathbb{R}P^{4}$ or the non-smoothable four-manifold $\ast\mathbb{R}P^{4}$ that is homotopy equivalent to the real projective four-space. We address the fourth step of our construction mechanism by computing the $\eta$-invariant [35] of the smooth four-manifolds involved in Section 6.4. Our main result in this case is the following theorem.
@@
Moreover, there is a non-smoothable closed topological four-manifold $\ast A(2p)$ that is homotopy equivalent to $A(2p)$.
@@ 1.1. Acknowledgements
V. B. is a member of GNSAGA – Istituto Nazionale di Alta Matematica ‘Francesco Severi’, Italy. R. T. thanks Stefan Bauer, Bob Gompf, Paul Kirk, Tye Lidman, Porter Morgan, Tom Mrowka, Mark Powell, Cliff Taubes, Mike Usher and Zoltan Szabó for useful e-mail exchanges.
@@ 2.1. Cut-and-paste procedure: oriented setting
where $r_{\theta}$ is the rotation of $S^{2}$ of angle $\theta$ around a fixed axis [17], [20, p. 156]. Such map is known as a Gluck twist. In order to prove Theorem 4, it is enough to show that the diffeomorphism $\varphi$ admits an extension $\Phi_{p}:S_{p}\rightarrow S_{p}$ for any odd $p\in\mathbb{Z}_{>0}$. We explicitly build the diffeomorphism $\Phi_{p}$ by using the identification
@@
coming from Lemma 2 as follows. Without loss of generality, we work in the setting of Remark 3. In particular, the deck transformation $\phi$ and the collection of three-balls $B_{1},\dots,B_{p}\subset D^{2}\times S^{1}$ are defined as in the proof of Lemma 2.
@@
where $r_{t}$ denotes the rotation of $S^{3}$ of an angle $2\pi t$ that restricts to a rotation along the $D^{2}$-factor on the solid torus $D^{2}\times S^{1}$ of the Heegaard decomposition of $S^{3}$ in Remark 3. Such solid torus contains the three-balls $B_{1},\dots,B_{p}$. Notice that for every $t\in[0,1]$, the map $F(t,\cdot)$ is an $f_{p}$-equivariant self-diffeomorphism of $S^{3}\setminus\sqcup_{i=1}^{p}B_{i}$, where $f_{p}$ can be chosen to act on the genus one Heegaard splitting of $S^{3}$ described in Remark 3. We can now define the automorphism
@@
As mentioned in the first step of our introduction, remove the tubular neighborhood of a framed simple loop $\alpha_{M}\subset M$ from a closed oriented four-manifold $M$ to obtain a compact oriented four-manifold $M\setminus\nu(\alpha_{M})$ with boundary $\partial(M\setminus\nu(\alpha_{M}))=S^{1}\times S^{2}$. Cap off its boundary with the compact four-manifold defined in (2.1) to construct a closed oriented four-manifold
@@
The reader is directed to [16, Chapter 5] regarding existence and uniqueness results of the tubular neighborhood of a locally flat simple loop in a topological four-manifold.
Independently of Corollary 5, the simplicity of the construction makes it straight-forward to compute several invariants of the new four-manifold (2.5) as we now sample.
@@
Let $M$ be a closed oriented four-manifold and let $\alpha_{M}\subset M$ be a simple framed loop whose homotopy class is $[\alpha_{M}]=0\in\pi_{1}(M)$. There is a diffeomorphism
@@
Lemma 2 and Lemma 7 have the following consequence.
Lemma 8.
Let $M(2)$ be a closed oriented four-manifold whose universal cover is $\widetilde{M}$ and let $\alpha\subset M(2)$ be a closed simple framed loop such that its homotopy class generates the group $\pi_{1}(M(2))=\mathbb{Z}/2$. If $p\in\mathbb{Z}_{>0}$ is odd, then the degree-two cover $\widehat{M}(p)$ of the closed oriented four-manifold
$M(2p)=(M(2)\setminus\nu(\alpha))\cup S_{p}$
is diffeomorphic to
$\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}.$
@@ 3.1. Raw materials for Theorem C
We will employ as a raw material the exotic four-manifolds $\mathcal{E}$ and $\mathcal{R}$ constructed in [29, Theorem 1.2].
@@
We refer the reader to [29, Section 6.2] for the precise definition of the four-manifolds $\mathcal{E}$ and $\mathcal{R}$. The four-manifolds of Theorem C are defined by applying the construction described in Section 2.1 to the exotic four-manifold $\mathcal{R}$ in the second item of Theorem 10.
@@
Alternatively, the existence of a homeomorphism as in Lemma 11 follows from work of Hambleton-Kreck [21, Theorem C]; cf. [34, Theorem 2.1].
@@
[Proof]This follows from the definition of $\mathcal{R}$ in [29] as the quotient of $\mathcal{E}$ by a fixed-point-free orientation-reversing involution, together with Lemma 8.
@@ 3.2. Diffeomorphism obstruction
In light of the definition of $\mathcal{E}$ in [29, Section 6.2], Proposition 13 implies the existence of a decomposition
@@
where $V=C_{0}\cup_{\phi_{0}\circ\sigma}Z$ is the building block defined in [29, Construction 6.4]. In particular, $C_{0}$ is a contractible four-manifold, $Z$ consists of five $-1$-framed $2$-handles attached to $\partial C_{0}$, $\partial V$ is the result of 00-surgery on $S^{3}$ along the Stevedore knot and $\zeta$ is any of its orientation-reversing fixed-point-free involutions. On the other hand, for every $p\geq 1$, $V_{p}$ is defined to be the four-manifold
@@
where $V^{\prime}=C_{0}\cup_{\phi_{0}}Z$ and $V^{\prime}_{p}=C_{0}^{p}\cup_{\phi_{0}}Z$ are made up by the same building blocks as $V$ and $V_{p}$ with different gluing maps. We refer the reader to [29, Construction 6.4] for further details.
@@
We proceed to provide some technical lemmas that are needed to distinguish the smooth structures of Theorem C. These results are generalizations of [29, Proposition 6.9] and [29, Lemma 6.3] respectively. In particular, we make use of the Ozsváth-Szabó four-manifold invariants defined in [31, Section 2.4].
In the following, we denote by $\Psi_{W,\mathfrak{t}}\in HF^{-}_{\text{red}}(\partial W,\mathfrak{t}|_{\partial W})$ the relative invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(W,\mathfrak{t})$ where $W$ is a compact four-manifold with non-empty boundary, and $\Phi_{X,\mathfrak{s},L}$ is the closed four-manifold invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s})$ with respect to a one-dimensional subspace $L\subset H_{2}(X;\mathbb{Q})$. We refer the reader to [29, Section 6.1] for the needed background on such invariants.
@@
[Proof]The case $p=1$ is [29, Proposition 6.9]. Since $C_{0}^{p}\setminus B^{4}$ is a rational homology cobordism between $S^{3}$ and $\partial C_{0}^{p}$, the same proof holds for $p>1$ by replacing $C_{0}$ with $C_{0}^{p}$ in the argument. The proof of the following result is a straightforward adaptation of the one of [29, Lemma 6.3].
@@
[Proof of Proposition 14] Item (1) follows from the construction and the fact that $\mathcal{E}^{\prime}$ is an exotic ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ by [29, Theorem 6.7]. By Lemma 7, there is a degree-two covering
@@
Let us now prove Item (2). In particular, $\mathcal{E}_{p}$ and $\mathcal{E}^{\prime}_{p}$ are homeomorphic by construction, since $\mathcal{E}$ and $\mathcal{E}^{\prime}$ are homeomorphic by [29, Theorem 6.7]. To conclude the proof, we need to show that the degree-two coverings $\mathcal{E}_{p}\rightarrow\mathcal{R}_{2p}$ and $\mathcal{E}^{\prime}_{p}\rightarrow\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ are not diffeomorphic. Consider the decomposition
@@
where $V_{a}$ and $V_{b}^{p}$ are copies of $V$ and $V_{p}$ respectively without the $T^{-}_{Q_{0}}$ component, see [29, Construction 6.4]. As for the case of $\mathcal{E}$, we see that $H_{2}(\mathcal{E}_{p})$ has a basis
@@
  • $(\Xi,\Theta)$ is the basis for $H_{2}(T_{Q_{0}})$ from [29, Remark 3.11]. In particular, $\Xi$ is represented by a capped-off Seifert surface for $Q_{0}$ in $S^{3}(Q_{0})$ and $\Theta$ is represented by a $2$-sphere of self-intersection $-2$ and $\Xi\cdot\Theta=1$.
As a consequence of Lemma 15 and 16, there is a ${\operatorname{Spin}}^{\mathbb{C}}$-structure $(\mathcal{E}_{p},\mathfrak{t})$ on such that $\Phi_{\mathcal{E}_{p},\mathfrak{t},L}\neq 0$, where $L=\text{Span}(\Xi)\subset H_{2}(\mathcal{E}_{p};\mathbb{Q})$. By [29, Lemma 6.2], we have that $\Xi$ cannot be represented by any surface of genus less than $2$. It is possible to show that if there was a diffeomorphism
@@
then one could build a spherical representative for $\Xi$ as in the proof of [29, Theorem 6.7], which would lead to a contradiction.